Tohban Report 2018-05-30
From RHESSI Wiki
|Start Date:||23 May 2018|
|End Date:||30 May 2018|
|Next Tohban:||Jim McTiernan|
|List all reports|
This week there were two C-class flares, one on the 23 and one on the 28th. At the end of last week, there were also a handful of B-class flares. Thus, we can say that solar activity was sort of high this week considering we are in quiet sun period. There is a B-type/beta sunspot group in NOAA 2712 which is currently in decay phase losing sunspot count and sunspot area. We had a spectacular active region at the center of the disk yesterday, matching the Hi-C sounding rocket flight and a Nustar campaign. The weak mixed polarities and weak magnetic shear located within the region's intermediate and trailing plage has simplified since the C2.7 event two days ago. We have solar activity in class B with the chance of another C-class event less likely in the next 24 hour period.
The SSR was over the 90% most of the week but currently is at 71%. It is now in a "stable" configuration.
All detectors are offline, we are still annealing.
Colt tips The cold tip 1 and 2 temperatures are ~93.4K and ~140.937K. and Cold plates are at ~166.742 and 165.365K.
Currently, the Cryocooler power is ~41.150 W and the acceleration is at ~21.966mg.
The general trend of the temperatures can be found here: http://sprg.ssl.berkeley.edu/~ayshih/soh/latest_temperatures.pdf
All detectors are off. We are still in annealing.
We are still annealing. We continue taking SAS data at high cadence this week. There was a set of IDLE problems due to errors in commanding during the weekend. We were in precesion mode what affected the spacecraft going to IDLE. The plan for this coming week is to change the parameter table that controls all these modes.
|HLAT Decimation||Rear decimation|
|Night time data (fronts)||plus/minus 4 minutes from daylight|
|Night time data (rears)||plus/minus 4 minutes from daylight|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Detector problems?||See notes above.|